A cat of mass 1 kg moves up  on an massless string which goes along a friction less pulley so as to just lift a block of mass 2 kg.After sometime the cat stops moving with respect to the string .The magnitude of change in cats acceleration is?

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Debadutta,

initial upward acceleration of cat=g

final acceleration=(2-1)g/(2+1)=g/3 upward

change in cat's acceleration=g-(g/3)=2g/3

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Debadutta !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR